A. For the above prices and quantities: Which are the demand quantities and which are the supply quantities?

1.            P            Q1              Q2

              500          1000            100

             1000          900             200

             1500          800             300

             2000          700             400

             2500          600             500

             3000          500             600

             3500          400             700

             4000          300             800

             4500          200             900

             5000          100            1000

     A. For the above prices and quantities: Which are the demand             

quantities and which are the supply quantities?

     B. Graph demand and supply on one graph. (Plot the points)

     C. What is the equilibrium price and quantity?

         Approximately.

      D. What would a price ceiling set at $4000 do to this market?

     E. What would a price ceiling set at $1000 do to this market?

      F. What would a price floor set at $4000 do to this market?

      G. Increase demand by 200% and answer a-f again.
Answer #1

(A)

The series Q1 shows that as price increases, quantity decreases, which follows the law of demand. So Q1 is the demand.

The series Q2 shows that as price increases, quantity increases, which follows the law of supply. So Q2 is the supply.

(B)

5000 ך 4500 4000 3500 -3000 U2500 2000 1500 1000 500 0 Demand -Supply 0 200 400 800 1000 QUANTITY600

(C)

Equilibrium is at intersection of Demand and Supply curves (point E). From graph,

Equilibrium price = $2750

Equilibrium quantity = 550

(D)

When price ceiling = $4000,

Quantity demanded (QD) = 300

Quantity supplied (QS) = 800

Since QD < QS, there is a surplus in the market.

Surplus = QS - QD = 800 - 300 = 500

NOTE: As per Answering Policy, 1st 4 parts are answered.